OpenStudy (anonymous):

What expression is the simplified form of sqrt y4?

OpenStudy (anonymous):

\[\sqrt{y4}\]

OpenStudy (anonymous):

\[y ^{2}\]

OpenStudy (anonymous):

thank you^^

OpenStudy (australopithecus):

If you want it explained \[\sqrt{y^{4}} = (y^{4})^{1/2}\] what is (1/2)*4 = 4/2 = 2 thus we have \[y^{2}\]

OpenStudy (australopithecus):

hope this helps you to understand radicals a bit better

OpenStudy (anonymous):

It does thank you